OpenStudy (anonymous):

Consider the basis S = {v1, v2, v3} for R3, where v1 = (1, 2, 1), v2 = (2, 9, 0), and v3 = (3, 3, 4), and let T:R3 → R2 be the linear operator such that T(v1) = (1, 0), T(v2) = (− 1, 1), T(v3) = (0, 1) Find a formula for T(x1, x2, x3), and use that formula to find T(9, 80, 9).

OpenStudy (perl):

so we have a transformation from a 3 vector to a 2 vector

OpenStudy (perl):

ok so far we have

OpenStudy (perl):

Consider the basis S = {v1, v2, v3} for R3, where v1 = (1, 2, 1), v2 = (2, 9, 0), and v3 = (3, 3, 4), and let T:R3 -> R2 be the linear operator such that T(v1) = (1, 0), T(v2) = (- 1, 1), T(v3) = (0, 1) Find a formula for T(x1, x2, x3), and use that formula to find T(9, 80, 9). Note i will use capital letters for bold letters We first express X = (x1,x2,x3) as a linear combination of V1 = (1,2,1) , V2 = (2,9,0) and V3 = (3,3,4). If we write (x1,x2,x3) = k1 ( 1,2,1) + k2( 2,9 ,0) + k3 (3,3,4) we obtain 1*k1 + 2*k2 + 3*k3 = x1 2*k1 + 9*k2 + 4*k3 = x2 1*k1+ 0*k2 + 4*k3 = x3

OpenStudy (perl):

now we do guassian elimination

OpenStudy (perl):

by the way, this is not the easiest way to solve this . this is the intro to the chapter im reading

OpenStudy (perl):

ok so let me get my maple computer

OpenStudy (perl):

err software,

OpenStudy (perl):

you still with me?

OpenStudy (perl):

check over for any errors

OpenStudy (anonymous):

Yes. yes. I'm following..

OpenStudy (perl):

and i already made a mistake :)

OpenStudy (perl):

should be 1*k1 + 2*k2 + 3*k3 = x1 2*k1 + 9*k2 + 3*k3 = x2 1*k1+ 0*k2 + 4*k3 = x3

OpenStudy (anonymous):

I kinda did something similar to your solution and i got (505, -124) but i'm not sure i did it right.. I happen to get the values of k1,k2, and k3 in terms of x1,x2, and x3..

OpenStudy (perl):

the reduced matrix is [[1,0,0, 21 x3-36 x1+8 x2], [0,1,0, -3 x3+5 x1-x2 ] [0,0,1, -5 x3+9 x1-2 x2]]

OpenStudy (perl):

k1 = 21x3 - 36x1 + 8x2 k2 =- 3 x3+5 x1-x2 k3 = -5 x3+9 x1-2 x2

OpenStudy (anonymous):

How did you get the reduced matrix? I got lost

OpenStudy (perl):

I used maple software. don't worry, we can go back to this lets see if this solution works (x1,x2,x3) = k1* ( 1,2,1) + k2*( 2,9 ,0) + k3*(3,3,4) substitute

OpenStudy (perl):

currently using maple 18 , thats what its called

OpenStudy (perl):

We found that k1 = 21x3 - 36x1 + 8x2 k2 =- 3 x3+5 x1-x2 k3 = -5 x3+9 x1-2 x2 We had before (x1,x2,x3) = k1* ( 1,2,1) + k2*( 2,9 ,0) + k3*(3,3,4) substitute (x1,x2,x3) = (21x3 - 36x1 + 8x2)* ( 1,2,1) +( - 3 x3+5 x1-x2 ) *( 2,9 ,0) + (-5 x3+9 x1-2 x2) *(3,3,4)

OpenStudy (perl):

T(x1,x2,x3) = (21x3 - 36x1 + 8x2) T(v1) + ( - 3 x3+5 x1-x2 )* T(v2) + (-5 x3+9 x1-2 x2) *T(v3)

OpenStudy (perl):

T(v1) = (1, 0), T(v2) =(1, 1), T(v3) = (0, 1) T(x1,x2,x3) = (21x3 - 36x1 + 8x2) *(1,0) + ( - 3 x3+5 x1-x2 )* (-1,1) + (-5 x3+9 x1-2 x2) *(0,1)

OpenStudy (perl):

T(x1,x2,x3) = ( (21x3 - 36x1 + 8x2) *(1)+ ( - 3 x3+5 x1-x2 )* (-1) + (-5 x3+9 x1-2 x2) *(0) , (21x3 - 36x1 + 8x2) *(0) + ( - 3 x3+5 x1-x2 )* (1) + (-5 x3+9 x1-2 x2) *(1))

OpenStudy (perl):

that simplifies to

OpenStudy (perl):

T(x1,x2,x3) = ( (21x3 - 36x1 + 8x2) *(1)+ ( - 3 x3+5 x1-x2 )* (-1) , ( - 3 x3+5 x1-x2 )* (1) + (-5 x3+9 x1-2 x2) *(1) )

OpenStudy (perl):

this simplifies to

OpenStudy (perl):

T(x1,x2,x3) = ( 21x3 - 36x1 + 8x2 + 3 x3-5 x1+x2 , - 3 x3+5 x1-x2 -5 x3+9 x1-2 x2 )

OpenStudy (anonymous):

505,-124?

OpenStudy (perl):

T(x1,x2,x3) = ( 24x3 - 41x1 + 9x2 , -8 x3+14 x1-3x2 )

OpenStudy (anonymous):

( - 3 x3+5 x1-x2 )* (-1) , ( - 3 x3+5 x1-x2 )* (1) <-- So this will not be cancelled. </3

OpenStudy (perl):

woops V1 = (1,2,1)

OpenStudy (perl):

lets test this show that T(V1) = (1,0) ie show that T ( 1,2,1) = (1,0)

OpenStudy (perl):

yes it works, just checked :)

OpenStudy (perl):

those are separated by commas, so they are in different places like ( x , y )

OpenStudy (perl):

T(9, 80, 9). plug into T(x1,x2,x3) = ( 24x3 - 41x1 + 9x2 , -8 x3+14 x1-3x2 )

OpenStudy (anonymous):

So it is 567, -186 :D

OpenStudy (perl):

correct

OpenStudy (anonymous):

Thank youu! <3

OpenStudy (perl):

ok lets streamline this approach

OpenStudy (perl):

or did you want to go back to see the gaussian elimination

OpenStudy (anonymous):

It's okay I guess. I followed your solution so i guess understand. :D

OpenStudy (perl):

1*k1 + 2*k2 + 3*k3 = x1 2*k1 + 9*k2 + 3*k3 = x2 1*k1+ 0*k2 + 4*k3 = x3 R2 ' = R2 + (-2)*R1 R3 ' = R3 + (-1)*R1

OpenStudy (perl):

i will be back later :) goodbye